# Bosonization in $R$-paraparticle Luttinger models

### v1

@@ 4 Potential experimental realization
The Luttinger model assumes a linearized single-particle dispersion which makes it valid only for particles with momentum close to $\pm k_{F}$. In the fermionic Luttinger model literature, all ideal one-dimensional systems exhibit spin-charge separation, but for realistic materials these conditions are not always met and systems that satisfy the above requirements are said to reside in the Luttinger liquid phase. Several works have experimentally verified the existence of the Luttinger liquid phase by observing some of its key signatures [47, 48, 49, 50, 51, 31]. We highlight a potential experimental design for observing paraparticle signatures by borrowing ideas from [31], which we outline as follows.
In the above paper, Luttinger liquid signatures were detected as two parabolic dispersions due to separate fermi seas with different masses, which are associated with spin and charge excitations. Since parafermionic Luttinger models also exhibit flavor-charge separation, we should expect that a one-dimensional system containing parafermions would also display such behavior. Current predictions show that paraparticles can occur as quasiparticle excitations [23, 8, 52, 53, 54]. In this case, the fermi momentum in one dimension, expressed as [55]
@@
is in general not equal for different parafermions. Here, $N_{0}$ denotes the number of particles occupying the ground state, and assuming that the system is made up of fundamental fermions which in turn have emergent parafermion quasiparticle properties, the number of fermions in the ground state is generally not equal to the number of parafermions in the quasiparticle ground state. By then constructing a system which does not exhibit spin-density excitations (e.g., there is no spin-charge separation or the system is fully spin-polarized) but contains emergent paraparticles that are in the Luttinger liquid phase, then signatures of flavor-charge separation may also be detected as separate dispersion relations as in [31].
The above proposal is roughly sketched, but it does give us a roadmap of research extensions to undertake in order to achieve an experimental observation of paraparticle signatures through Luttinger liquids. Firstly, a method which can turn a system of ordinary fermions or bosons to quasiparticle parafermion excitation is required. In [23], it was shown that emergent one-dimensional quasiparticle excitations with parastatistical GCRs can be developed from a generic spin model where the spin operators have $R$-paraparticle behaviors. For example, emergent paraparticles described by Eq. (13a) were shown to be realizable using three-level Rydberg atom or molecule systems [56, 57], while those described by Eq. (13b) with $\alpha=1$ can be modelled using systems described in [58]. In general, several papers have proposed a method of mapping between particles of different statistics in one-dimensional systems [59, 60], although an extension to the $R$-paraparticles is yet to be explored.
@@
Further complications to this proposal would be due to the limitations of our results to paraparticles with $2$ internal degrees of freedom. The construction of the flavor wave operator was inspired using an analog of spin waves in spin-$\frac{1}{2}$ systems. While quantum spins are not directly observable, electron spin-wave propagation can be observed experimentally [42]. As such, Eq. (15d) should be sufficient to describe a wave propagation due to some internal degree of freedom of a paraparticle with $m=2$. This allows us to potentially observe paraparticle signatures even without full knowledge of the internal symmetries at play within the system. However, if the emergent paraparticle contains $m>2$ internal degrees of freedom, then the flavor wave operator must be redefined to account for multiple types of flavor wave oscillations, while at the same time, preserving the total degrees of freedom of the system.
@@ 5 Summary and Discussion
We have shown that bosonization applies to certain one-dimensional $R$-paraparticles systems described by the Luttinger model, to which flavor-charge separation occurs as a result. These paraparticles are classified as $R$-parafermions, which obey general exclusion principles and thus can only occupy a limited number of particle states. Parafermionic density waves are shown to have bosonic signatures, but flavor waves generally do not exhibit this behavior. A comparison of the partition function in the Luttinger model shows that only parafermions that satisfy Pauli’s exclusion principle preserve their degeneracy when changing from the parafermion basis to the boson basis, further limiting the applicability of bosonization to paraparticles of order $p=1$. Nonetheless, these results show promises in potentially observing paraparticle signatures in one-dimensional systems.
We have motivated this study as a means of finding a general means of solving a paraparticle system in the presence of interaction. A dedicated section for such method was omitted in this paper since, by showing that a collection of parafermion operators can be recast as a boson operator, bosonization automatically implies that solutions applicable to fermionic Luttinger models should also apply to parafermionic systems (see Sec. 4.4C of [29]). Nonetheless, other generalizations such as the field-theoretic formalism of the Luttinger model should not be treated as trivial given the extensive use of generalized commutation relations in their definitions.
As mentioned in the previous section, generalization to arbitrary $m$ cases would benefit the development of possible experimental setups that may result to observations of paraparticle systems. This goes hand-in-hand with an understanding of the internal symmetries at play within the paraparticle. At the same time, a physical interpretation of a non-bosonic flavor wave oscillation is lacking, which may or may not affect potential observations of paraparticle singatures. Speculatively, this could also mean that $\hat{\sigma}_{r}(q)$ generally possess parabose statistics, but this remains an assumption until an example with such properties can be presented.
@@ Appendix A R-parabosons
In Sec. 2.2, we discussed the behavior of paraparticles where $d_{n>n^{\prime}}=0$ for some finite $n^{\prime}$. We now discuss what happens in the limit $n^{\prime}\to\infty$. We note that the single-mode partition functions of paraparticles satisfy [23]
@@
  • By definition, $z_{R}(-x)=1+\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}$ ($d_{0}=0$ for all paraparticles, i.e., there is always only one vacuum state), so the second equality in Eq. (A.2) requires $1>\allowbreak|\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}|$. $1>-\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}$ is automatically satisfied by the fact that $z_{-R}(x)>\allowbreak 0$ as required from a partition function.
@@
hence at the domain $\epsilon\in(\epsilon_{0},\infty)$, the average occupation number is exponentially decreasing from infinity, reminiscent of bose statistics. This motivates us to classify paraparticles that have single-mode partition functions in the form of Eq. (A.2) with $D_{n}\neq 0$ for all $n$ as $R$-parabosons.
@@
will always diverge at the radius of convergence $x_{0}$ of $z(x)$. This is because $xz^{\prime}(x)$ grows faster than $z(x)$ as $x\to x_{0}^{-}$, and thus the resulting behavior of $\langle\tilde{n}\rangle$ is $R$-parabosonic. An implication of this result is that in the $R$-paraparticle formulation, there are more parabosons than parafermions.

### v2

@@ 4 Potential experimental realization
The Luttinger model assumes a linearized single-particle dispersion which makes it valid only for particles with momentum close to $\pm k_{F}$. In the fermionic Luttinger model literature, all ideal one-dimensional systems exhibit spin-charge separation, but for realistic materials these conditions are not always met and systems that satisfy the above requirements are said to reside in the Luttinger liquid phase. Several works have experimentally verified the existence of the Luttinger liquid phase by observing some of its key signatures [47, 48, 49, 50, 51, 31]. We highlight a potential experimental design for observing $R$-paraparticle signatures by borrowing ideas from [31], which we outline as follows.
In the above paper, Luttinger liquid signatures were detected as two parabolic dispersions due to separate fermi seas with different masses, which are associated with spin and charge excitations. Since $R$-parafermionic Luttinger models also exhibit flavor-charge separation, we should expect that a one-dimensional system containing $R$-parafermions would also display such behavior. Current predictions show that $R$-paraparticles can occur as quasiparticle excitations [23]. In this case, the fermi momentum in one dimension, expressed as [55]
@@
is in general not equal for different $R$-parafermions. Here, $N_{0}$ denotes the number of particles occupying the ground state, and assuming that the system is made up of fundamental fermions which in turn have emergent $R$-parafermion quasiparticle properties, the number of fermions in the ground state is generally not equal to the number of $R$-parafermions in the quasiparticle ground state. By then constructing a system which does not exhibit spin-density excitations (e.g., there is no spin-charge separation or the system is fully spin-polarized) but contains emergent $R$-paraparticles that are in the Luttinger liquid phase, then signatures of flavor-charge separation may also be detected as separate dispersion relations as in [31].
The above proposal is roughly sketched, but it does give us a roadmap of research extensions to undertake in order to achieve an experimental observation of $R$-paraparticle signatures through Luttinger liquids. Firstly, a method which can turn a system of ordinary fermions or bosons to quasiparticle $R$-parafermion excitation is required. In [23], it was shown that emergent one-dimensional quasiparticle excitations with parastatistical GCRs can be developed from a generic spin model where the spin operators have $R$-paraparticle behaviors. For example, emergent $R$-paraparticles described by Eq. (13a) were shown to be realizable using three-level Rydberg atom or molecule systems [56, 57], while those described by Eq. (13b) with $\alpha=1$ can be modelled using systems described in [58]. In general, several papers have proposed a method of mapping between particles of different statistics in one-dimensional systems [59, 60], although an extension to the $R$-paraparticles is yet to be explored.
@@
Further complications to this proposal would be due to the limitations of our results to $R$-paraparticles with $2$ internal degrees of freedom. The construction of the flavor wave operator was inspired using an analog of spin waves in spin-$\frac{1}{2}$ systems. While quantum spins are not directly observable, electron spin-wave propagation can be observed experimentally [42]. As such, Eq. (15d) should be sufficient to describe a wave propagation due to some internal degree of freedom of an $R$-paraparticle with $m=2$. This allows us to potentially observe $R$-paraparticle signatures even without full knowledge of the internal symmetries at play within the system. However, if the emergent $R$-paraparticle contains $m>2$ internal degrees of freedom, then the flavor wave operator must be redefined to account for multiple types of flavor wave oscillations, while at the same time, preserving the total degrees of freedom of the system.
@@ 5 Summary and Discussion
We have shown that bosonization applies to certain one-dimensional $R$-paraparticles systems described by the Luttinger model, to which flavor-charge separation occurs as a result. These $R$-paraparticles are classified as $R$-parafermions, which obey general exclusion principles and thus can only occupy a limited number of particle states. $R$-parafermionic density waves are shown to have bosonic signatures, but flavor waves generally do not exhibit this behavior. A comparison of the partition function in the Luttinger model shows that only $R$-parafermions that satisfy Pauli’s exclusion principle preserve their degeneracy when changing from the $R$-parafermion basis to the boson basis, further limiting the applicability of bosonization to $R$-paraparticles of order $p=1$. Nonetheless, these results show promises in potentially observing $R$-paraparticle signatures in one-dimensional systems.
We have motivated this study as a means of finding a general means of solving an $R$-paraparticle system in the presence of interaction. A dedicated section for such method was omitted in this paper since, by showing that a collection of $R$-parafermion operators can be recast as a boson operator, bosonization automatically implies that solutions applicable to fermionic Luttinger models should also apply to $R$-parafermionic systems (see Sec. 4.4C of [29]). Nonetheless, other generalizations such as the field-theoretic formalism of the Luttinger model should not be treated as trivial given the extensive use of generalized commutation relations in their definitions.
As mentioned in the previous section, generalization to arbitrary $m$ cases would benefit the development of possible experimental setups that may result to observations of $R$-paraparticle systems. This goes hand-in-hand with an understanding of the internal symmetries at play within the $R$-paraparticle. At the same time, a physical interpretation of a non-bosonic flavor wave oscillation is lacking, which may or may not affect potential observations of $R$-paraparticle singatures. Speculatively, this could also mean that $\hat{\sigma}_{r}(q)$ generally possess parabose statistics, but this remains an assumption until an example with such properties can be presented.
@@ Appendix A R-parabosons
In Sec. 2.2, we discussed the behavior of $R$-paraparticles where $d_{n>n^{\prime}}=0$ for some finite $n^{\prime}$. We now discuss what happens in the limit $n^{\prime}\to\infty$. We note that the single-mode partition functions of $R$-paraparticles satisfy [23]
@@
  • By definition, $z_{R}(-x)=1+\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}$ ($d_{0}=0$ for all $R$-paraparticles, i.e., there is always only one vacuum state), so the second equality in Eq. (A.2) requires $1>\allowbreak|\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}|$. $1>-\sum_{n=1}^{n^{\prime}}d_{n}(-x)^{n}$ is automatically satisfied by the fact that $z_{-R}(x)>\allowbreak 0$ as required from a partition function.
@@
hence at the domain $\epsilon\in(\epsilon_{0},\infty)$, the average occupation number is exponentially decreasing from infinity, reminiscent of bose statistics. This motivates us to classify $R$-paraparticles that have single-mode partition functions in the form of Eq. (A.2) with $D_{n}\neq 0$ for all $n$ as $R$-parabosons.
@@
will always diverge at the radius of convergence $x_{0}$ of $z(x)$. This is because $xz^{\prime}(x)$ grows faster than $z(x)$ as $x\to x_{0}^{-}$, and thus the resulting behavior of $\langle\tilde{n}\rangle$ is $R$-parabosonic. An implication of this result is that in the $R$-paraparticle formulation, there are more $R$-parabosons than $R$-parafermions.